OpenStudy (anonymous):

34 + 12 ÷ 3 - (1-9) Help me simplify please.

OpenStudy (anonymous):

you need to do the division first, and the stuff in parentheses... so 12/3 = 4 and 1-9 = -8 then 34 + 4 - (-8) two negatives make a positive... so 34 + 4 +8 =??

OpenStudy (lilg132):

oops

OpenStudy (anonymous):

46 cannot be the answer it is not on the options given

OpenStudy (anonymous):

did you copy the problem exactly?